At 4 pm, Gigi and Madel left the music house. Gigi walked east and reached her house at 5 pm. Madel rode her bike west at 4:15 pm. Madel traveled twice as fast as Gigi. Their houses are 9 kilometers apart. How far did each travel?
Answer by stanbon(75887)

At 4 pm, Gigi and Madel left the music house. Gigi walked east and reached her house at 5 pm. Madel rode her bike west at 4:15 pm. Madel traveled twice as fast as Gigi. Their houses are 9 kilometers apart. How far did each travel?
-------------
Gigi DATA:
Time = 1 hr ; Distance = x km ; Rate = d/t = x kph
------------
Madel DATA:
Distance = 9-x km ; Rate = 2x kph ; Time = (9-x)/2x hrs
----------------
Comment: If Madel reached her house at the same time as Gigi reached
her house Madel's time is (3/4)hr.
---------------
EQUATION:
(9-x)/2x = 3/4
6x = 36-4x
10x = 36
x = 3.6 km (Gigi traveled this far)
9-3.6 = 5.4 (Mdel traveled this far)
==============
Cheers,
Stan H.